Evaluate the function f(x) =x² +7x-7 at the given values of the independent variable and simplify
a. F(6) b. f(x +4) c. f(- x)

Answer by alegriajoy33(14)

a) you plug in 6 for every X variable.
so your equation will look like. (6^2)+(7*6)-7 therefore your answer for A is 71
b) you plug in (x+4) for every X variable.
your equation looks like. ((x+4)^2)+7(x+4)-7 now you foil the (x+4)^2 so that is (x+4)(x+4) now it is F.O.I.L. F=First so that is the first number in the first set of perentheses and the first in the second. therefore it is X*X which is x^2. O=outer this would be the first in first set of perenthesese and the second in the second set. So this is X*4 therefore you get 4X. I=Inner so that is the second in the first set of perenthesese and first in second set. so it is 4*X therefore it is 4X. L=last this is the second in the first set of perenthesese times the first in the second set. so this is 4*4 which is 16.
s your equation so far is (X^2)+4x+4x+16 now combine like terms you get
(x^2)+8x+16 okay now you gotta do the second part of this problem and distribute the 7 t the (x+4) so multiply 7 by each term. you get 7x+28 so combine this with what you had before and the -7 so you should get in the end for the answer to b. (x^2)+15x+37.
C) Plug in (-X) in for x (-x^2)+7(-x)-7 now simplify it so you get (-X^2)-7X-7
so there you go. I hope you understand this. i is kinda hard to explain over
e-mail. so here you go tehehe. :0)